A resistance R and inductance L=0.01 H and a capacitance are connected in series when a voltage V=400Cos (300t-10) v is applied to the series combination, the current flowing I=10 √2cos (300t-55°)A. Find P and C.
1
Expert's answer
2019-03-21T16:16:52-0400

The active power:


"P=\frac{1}{2}VI\text{cos}(\phi_V-\phi_I)=""=\frac{1}{2}\cdot400\cdot10\sqrt{2}\text{cos}(-10^\circ-(-55^\circ))=2000\text{ W}."

Find the resistance:


"R=\frac{P}{I^2}=\frac{2000}{10^2}=20\space\Omega."

Write Ohm's law for the circuit:


"Z=\sqrt{R^2+(\omega L-1/\omega C)^2}=\frac{V}{I},"

"20^2+(300\cdot0.01-\frac{1}{300 C})^2=(\frac{400}{10\sqrt{2}})^2,"

"C=144.93\space\mu\text{F}"